# Map-level baryonification: Efficient modelling of higher-order correlations in the weak lensing and thermal Sunyaev-Zeldovich fields

### v1

@@ 4.1.3 Dependence on ellipticity
It is therefore important to determine the impact of the ellipticity correlations in the baryonification pipeline, for modelling the statistics of interest to us. The implementation of ellipticities in our model is trivial. The halo-centric distance to a pixel is now scaled as $x^{2}+y^{2}\rightarrow x^{2}+y^{2}/q^{2}$ where $q\leq 1$ is the ratio of the minor axis to the major axis. [^14] We only explore the impact of ellipticity on the 2D field, where $q$ is the only relevant shape parameter. Under this formalism, the major axis is assumed to be of length $R_{\rm 200c}$. In practice, this scaling is done by transforming the coordinate system — rotating the pixel grid so the major/minor axis of the halo is along the x-axis and y-axis of the grid — and then rescaling the $y$ distance by $q$ as denoted above. The relevant quantity, a displacement or gas pressure, is computed at the location given by this rescaled distance. We perform this test with IllustrisTNG, where we computed the orientations of halos using the same pipeline as Anbajagane et al. (2022a); this work computes the mass intertia tensor in 3D to compute the 3D ellipticity. In our case, we take only the 2$\times$2 subset of the tensor corresponding to the x and y directions, and compute a 2D ellipticity using the new tensor.
@@ 4.3 Mass dependence
In Section 4.2 we show that our model is flexible enough to jointly fit the simulation measurements within their statistical uncertainties. We now use this model to check the mass-dependence of the predicted signal for the different moments. This is analogous to similar plots made for the two-point function in Pandey et al. (2022, see their Figure 2) or To et al. (2024, see their Figure 3), and lend insight into the origin of the signal in different measurements. In general, we expect higher-order moments to be more sensitive to more massive halos than the 2nd moments, and similarly, for the tSZ field to be more sensitive to more massive halos than the lensing field.
@@ 5 Forecasting
The lensing field is modelled using the same pipeline from Anbajagane et al. (2023, 2024b). The maps are generated from the density fields of the Ulagam simulations, and then processed to contain all the observational effects found in the data, including the relevant survey mask/area and source galaxy redshift distribution. The two lensing surveys we focus on are: (i) the Dark Energy Survey (DES, The Dark Energy Survey Collaboration 2005), which is an optical imaging survey of 5,000 deg2 of the southern sky, and is currently the largest precision photometric dataset for cosmology. The Year 3 data products and cosmology results are available (Sevilla-Noarbe et al. 2021; Abbott et al. 2022), while the legacy Year 6 dataset is not yet available at the time of publishing this work. We focus on this Year 6 dataset. We then consider (ii) the Rubin Observatory Legacy Survey of Space and Time (LSST), which is a 14,000 deg2 survey that probes higher redshifts, and is the successor to current weak lensing surveys. We forecast only for the final, year 10 dataset. The characteristics of the surveys (shape noise, and $n(z)$ distribution parameters) are given in Table 3. More details can be found in Appendix A.1
@@
The inclusion of higher-order moments (blue and yellow contours) greatly improves the constraints over using the 2nd moments alone (purple contours). This is expected as baryonic imprints are tied to the astrophysical processes within halos and are therefore stronger on non-linear scales. These scales are also where the field is more non-Gaussian, and parameter constraints improve from including measurements of any higher-order correlations at these scales. Aricò et al. (2021a) already show that a range of baryonification parameter choices can accurately fit a matter power spectrum from simulations but only a subset of those can also jointly fit the matter bispectrum, indicating that the addition of higher-order information constrains the parameter space further. We expand on this by using up to 4th order in the fields, and also by using both the lensing and tSZ fields. Figure 10 also shows the marginalized constraints from using only the 3rd moments (4th moments), presented as a green (peach) dashed line in the 1D posteriors along the diagonal panels of the triangle plot. [^15] For the 4th moments, we have used only the connected piece, $\langle x_{1}x_{2}x_{3}x_{4}\rangle-(\langle x_{1}x_{2}\rangle\langle x_{3}x_{4}\rangle+\text{perms.})$. ), so that there is a clean separation between the information in the 2nd moments and 4th moments. In many cases, these higher order moments are more constraining than the 2nd moments alone.
@@
Figure 11: Fisher constraints on the baryonification and cosmological parameters from different combinations of the convergence and tSZ field as measured by DES Y6 and SPT 3G. The datavector uses the 2nd and 3rd moments. The convergence field and tSZ field have different sensitivities to the parameters, and the degeneracy directions are often orthogonal. As a result the combination of fields is significantly more constraining.
@@
The field-by-field comparison of Figure 11 also highlights the improvement in cosmology constraints ($\Omega_{\rm m}$ and $\sigma_{8}$) due to the addition of the tSZ field. The cross-correlation of the lensing and tSZ fields is a significant factor in this improvement. The results also indicate that the tSZ-only analysis is unable to jointly constrain baryonification models alongside cosmology. However, this result should not be over-interpreted or generalized since our analysis explicitly limits the tSZ measurements to $\arcmin\theta>3\arcmin$, whereas the tSZ power spectrum has the highest signal-to-noise at even smaller scales, $\ell\sim 3000$ (Raghunathan et al. 2022b, see their Figure 4). Our tSZ maps are generated at $\texttt{NSIDE}=1024$ — the same choice as that of the lensing field — whereas the maps from current SPT datasets are made at $\texttt{NSIDE}=8192$ (Bleem et al. 2022). Thus, there is more usable information in the tSZ field that we have not accessed in our datavectors given we limit our maps (and therefore our analysis) to a coarser resolution. This is done primarily in the interest of minimizing the computation cost, as generating a tSZ map of $\texttt{NSIDE}=8192$ would take around 64×64$\times$64 × longer than generating one of $\texttt{NSIDE}=1024$. Note that for the lensing analysis, the choice of $\texttt{NSIDE}=1024$ is realistic.
@@
All analyses above, summarised in Table 2, show that a number of analysis setups result in parameter uncertainties that are of the same order as the fiducial parameter values listed in Table 1: (i) the $\nu_{X}$ parameters, which characterize the redshift evolution of $\theta_{\rm ej}$ and $M_{\rm c}$, (ii) $\gamma$ which is the gas slope at $R\sim R_{\rm ej}$, and (iii) the mass-dependence parameters $\mu_{X}$. This may indicate that robust predictions of the lensing and tSZ moments can be obtained while fixing these parameters. However, a robust validation of this statement requires fitting this model (with and without varying the parameters mentioned here) to mock surveys derived from different hydrodynamical simulations. Appendix D also presents results that indicate the $\nu_{X}$ parameters may not be necessary for jointly modelling the moments measurements across redshift.
@@ 6 Conclusions
The baryonification method has rapidly become a important tool in characterizing the two-point correlation functions measured by lensing surveys. The method, however, is a general prescription on how to add baryonic fields to N-body simulations through an approximate, phenomological approach. In this work, we use this formalism to model the lensing and tSZ field. One could similarly use this prescription to model other fields that are observables of gas physics, such as the kinematic SZ effect, which has already been measured using different combinations of CMB and galaxy surveys (e.g., Mallaby-Kay et al. 2023; Hadzhiyska et al. 2024), and the X-ray luminosity, measured for a large fraction of the sky by the eROSITA survey (Bulbul et al. 2024). In fact, some of these observables have already been used to constrain baryonification models (Schneider et al. 2019, 2022; Grandis et al. 2024). Thus, we have only probed some aspects of the full potential provided by the baryonification formalism.
@@ Acknowledgements
We thank Marco Gatti and Yuuki Omori for discussions during early stages of this work, Srini Raghunathan for guidance on the tSZ forward models used in Raghunathan et al. (2022a, b), and Kayla Kornoelje for details on the linear combination method used by SPT. The conception and pursuit of this work arose from many conversations between DA and Eric Baxter. The authors are grateful to Eric for his collaboration and friendship over the many years, and for his efforts in making our research communities vibrant avenues for discussion and discovery. He will be missed.
@@
The baryonification pipeline is publicly available at https://github.com/DhayaaAnbajagane/Baryonification.
@@ A.1 Weak lensing
Survey mask & mass map construction. The noisy shear field — which is the sum of the true shear fields and the shape noise fields — is then masked according to the survey footprint. For DES Y6 we use the provided survey mask in the year 3 data release (Sevilla-Noarbe et al. 2021). For LSST Y1/Y10, we divide the sky into three equal-area cutouts of roughly 14,000 deg2 each, which is the expected area coverage for Y10 and $\approx 15\%$ larger than the expected area for Y1. The noisy shear maps are converted to convergence using the relation in Equation 41. We only use the resulting E-mode field, $\kappa_{E}$, for our analyses. This follows the same procedures used in Chang et al. (2018); Jeffrey et al. (2021).
@@ C.2 Sensitivity to non-thermal pressure models
Figure 16 compares this model to other alternatives, particularly the predictions of Green et al. (2020). The latter model depends only on the peak height of the halo, and is therefore a prediction and not a fit to the data. We find that our model parameterization is flexible enough to capture features of the form predicted by Green et al. (2020), which in turn is a advanced version of previous non-thermal models, such as that of Nelson et al. (2014).
@@ Appendix D Validation at higher redshifts
Appendix E Computational details

### v2

@@ 4.1.3 Dependence on ellipticity
It is therefore important to determine the impact of the ellipticity correlations in the baryonification pipeline, for modelling the statistics of interest to us. The implementation of ellipticities in our model is trivial. The halo-centric distance to a pixel is now scaled as $x^{2}+y^{2}\rightarrow x^{2}+y^{2}/q^{2}$ where $q\leq 1$ is the ratio of the minor axis to the major axis. [^17] We only explore the impact of ellipticity on the 2D field, where $q$ is the only relevant shape parameter. Under this formalism, the major axis is assumed to be of length $R_{\rm 200c}$. In practice, this scaling is done by transforming the coordinate system — rotating the pixel grid so the major/minor axis of the halo is along the x-axis and y-axis of the grid — and then rescaling the $y$ distance by $q$ as denoted above. The relevant quantity, a displacement or gas pressure, is computed at the location given by this rescaled distance. We perform this test with IllustrisTNG, where we computed the orientations of halos using the same pipeline as Anbajagane et al. (2022a); this work computes the mass inertia tensor in 3D to compute the 3D ellipticity. In our case, we take only the 2$\times$2 subset of the tensor corresponding to the x and y directions, and compute a 2D ellipticity using the new tensor.
@@ 4.2 Comparison to hydro simulation
Under the current statistical uncertainties, we find that our model is able to jointly fit the simulation data at a fractional accuracy that is at worst between $5\%$ to $\approx 50\%$, depending on the exact moment being considered, with the fractional difference increasing for moments with larger measurement uncertainties (for example, $\langle y^{3}\rangle$ or $\langle y^{4}\rangle$, whose signal is dominated by just a few halos in the TNG volume). Given the caveats related to TNG that we have discussed throughout the text above, a more robust analysis is needed to calibrate the baryonification model accuracy to much better precision. Such an analysis would require a much larger simulation; for example Bahamas (McCarthy et al. 2017) or Millenium-Tng (Pakmor et al. 2023). We leave these efforts to future work.
@@ 4.3 Mass dependence
In Section 4.2 we show that our model is flexible enough to jointly fit the simulation measurements within their statistical uncertainties. We now use this model to check the mass-dependence of the predicted signal for the different moments. This is analogous to similar plots made for the two-point function in Pandey et al. (2022, see their Figure 2) or To et al. (2024, see their Figure 3), and lend insight into the origin of the signal in different measurements. In general, we expect higher-order moments to be more sensitive to more massive halos than the 2nd moments, and similarly, for the tSZ field to be more sensitive to more massive halos than the lensing field. The former is because taking higher powers of a variable amplify the tails of its distribution, and the density field has a strongly positive-skewed tail — occupied by massive halos — whose presence/importance is amplified by taking higher powers of the field. The latter is because weak lensing depends linearly on the density field and therefore mass, whereas the tSZ signal follows $\propto M^{5/3}$ (e.g., Lee et al. 2022, see their Figure 1 for a comparison across simulations) and therefore is sensitive to higher masses than lensing.
@@ 5 Forecasting
The lensing field is modelled using the same pipeline from Anbajagane et al. (2023, 2024b). The maps are generated from the density fields of the Ulagam simulations (Section 3.3), and then processed to contain all the observational effects found in the data, including the relevant survey mask/area and source galaxy redshift distribution. The two lensing surveys we focus on are: (i) the Dark Energy Survey (DES, The Dark Energy Survey Collaboration 2005), which is an optical imaging survey of 5,000 deg2 of the southern sky, and is currently the largest precision photometric dataset for cosmology. The Year 3 data products and cosmology results are available (Sevilla-Noarbe et al. 2021; Abbott et al. 2022), while the legacy Year 6 dataset is not yet available at the time of publishing this work. We focus on this Year 6 dataset. We then consider (ii) the Rubin Observatory Legacy Survey of Space and Time (LSST), which is a 14,000 deg2 survey that probes higher redshifts, and is the successor to current weak lensing surveys. We forecast only for the final, year 10 dataset. The characteristics of the surveys (shape noise, and $n(z)$ distribution parameters) are given in Table 3. More details can be found in Appendix A.1
@@
We note that all our analyses use relatively “medium”-resolution data-products: the Ulagam simulations are run with $512^{3}$ particles, and produce maps of $\texttt{NSIDE}=1024$. These choices set a limit on the scales we can analyse. Simulations of higher resolution enable access to even smaller scales, which would place tighter constraints on the baryonic processes discussed in this work. Therefore all our constraints below could be improved further by pushing the analysis to even smaller scales. In this work, we are limited by the resolution of the simulations available for a Fisher forecast.
These results should also not be over-interpreted or generalized since our analysis explicitly limits the tSZ measurements to $\arcmin\theta>3\arcmin$, whereas the tSZ power spectrum has the highest signal-to-noise at even smaller scales, $\ell\sim 3000$ (Raghunathan et al. 2022b, see their Figure 4). Our tSZ maps are generated at $\texttt{NSIDE}=1024$ — the same choice as that of the lensing field — whereas the maps from current SPT datasets are made at $\texttt{NSIDE}=8192$ (Bleem et al. 2022). Thus, there is more usable information in the tSZ field that we have not accessed in our datavectors given we limit our maps (and therefore our analysis) to a coarser resolution. This is done primarily in the interest of minimizing the computation cost, as generating a tSZ map of $\texttt{NSIDE}=8192$ would take around 64×64$\times$64 × longer than generating one of $\texttt{NSIDE}=1024$. Note that for the lensing analysis, the choice of $\texttt{NSIDE}=1024$ is realistic, though it may be possible to access smaller scales ($\texttt{NSIDE}=2048,4096$) — if the lensing systematics are better controlled/modelled — which would improve our constraints. The constraints from both fields could also improve/deteriorate depending on the summary statistics being considered; for examples in the lensing field, see Euclid Collaboration et al. (2023); Anbajagane et al. (2023); Gatti et al. (2024a). We have not explored this direction, and have limited ourselves to the moments in this work, as they have already been used to constrain cosmology (Gatti et al. 2022, 2024a).
@@
The inclusion of higher-order moments (blue and yellow contours) greatly improves the constraints over using the 2nd moments alone (purple contours). This is expected as baryonic imprints are tied to the astrophysical processes within halos and are therefore stronger on non-linear scales. These scales are also where the field is more non-Gaussian, and parameter constraints improve from including measurements of any higher-order correlations at these scales. Aricò et al. (2021a) already show that a range of baryonification parameter choices can accurately fit a matter power spectrum from simulations but only a subset of those can also jointly fit the matter bispectrum, indicating that the addition of higher-order information constrains the parameter space further. We expand on this by using up to 4th order in the fields, and also by using both the lensing and tSZ fields. Figure 10 also shows the marginalized constraints from using only the 3rd moments (4th moments), presented as a green (peach) dashed line in the 1D posteriors along the diagonal panels of the triangle plot. [^18] For the 4th moments, we have used only the connected piece, $\langle x_{1}x_{2}x_{3}x_{4}\rangle-(\langle x_{1}x_{2}\rangle\langle x_{3}x_{4}\rangle+\text{perms.})$. ), so that there is a clean separation between the information in the 2nd moments and 4th moments. In many cases, the “only 3rd moments” and “only 4th moments” constraints are more constraining than the 2nd moments alone (though this is not the case for the cosmology parameters).
@@
Figure 11: Fisher constraints on the baryonification and cosmological parameters from different combinations of the convergence and tSZ field as measured by DES Y6 and SPT 3G. The datavector uses the 2nd and 3rd moments. The convergence field and tSZ field have different sensitivities to the parameters, and the degeneracy directions are often orthogonal. As a result the combination of fields is significantly more constraining. The 1D marginal distributions of $\Omega_{m}$ and $\sigma_{8}$ also show a dotted blue line, which is the lensing-only constraints without marginalizing over any baryonification parameters.
@@
The field-by-field comparison of Figure 11 also highlights the improvement in cosmology constraints ($\Omega_{\rm m}$ and $\sigma_{8}$) due to the addition of the tSZ field. The cross-correlation of the lensing and tSZ fields is a significant factor in this improvement. The results also indicate that the tSZ-only analysis is unable to jointly constrain baryonification models alongside cosmology. However, this result should not be over-interpreted or generalized since our analysis explicitly limits the tSZ measurements to $\arcmin\theta>3\arcmin$, whereas the tSZ power spectrum has the highest signal-to-noise at even smaller scales, $\ell\sim 3000$ (Raghunathan et al. 2022b, see their Figure 4). Our tSZ maps are generated at $\texttt{NSIDE}=1024$ — the same choice as that of the lensing field — whereas the maps from current SPT datasets are made at $\texttt{NSIDE}=8192$ (Bleem et al. 2022). Thus, there is more usable information in the tSZ field that we have not accessed in our datavectors given we limit our maps (and therefore our analysis) to a coarser resolution. This is done primarily in the interest of minimizing the computation cost, as generating a tSZ map of $\texttt{NSIDE}=8192$ would take around 64×64$\times$64 × longer than generating one of $\texttt{NSIDE}=1024$. For the lensing analysis, the choice of $\texttt{NSIDE}=1024$ is realistic, with some potential to extend to higher resolution if the systematics can be controlled robustly.
The 1D marginal distributions in Figure 11 also shows a dotted blue line, which are the results from a lensing-only analysis (using 2nd and 3rd moments) but with all baryonification parameters fixed. The constraints on cosmology — $\sigma(\Omega_{m})=0.008$ and $\sigma(\sigma_{8})=0.013$ — are comparable to those from using all 2nd and 3rd auto/cross-moments of the lensing and tSZ field after marginalizing over all baryonification parameters. Note that this compares lensing-only with a lensing plus tSZ analysis. In Appendix E, we consistently use a lensing plus tSZ analysis, with all baryonification parameters either fixed or varied in different stages, and find the constraints on cosmology degrades by $\approx 60\%$ if we vary all baryonification parameters.
@@
All analyses above, summarised in Table 2, show that a number of analysis setups result in parameter uncertainties that are of the same order as the fiducial parameter values listed in Table 1: (i) the $\nu_{X}$ parameters, which characterize the redshift evolution of $\theta_{\rm ej}$ and $M_{\rm c}$, (ii) $\gamma$ which is the gas slope at $R\sim R_{\rm ej}$, and (iii) the mass-dependence parameters $\mu_{X}$. This may indicate that robust predictions of the lensing and tSZ moments can be obtained while fixing these parameters. However, a robust validation of this statement requires fitting this model (with and without varying the parameters mentioned here) to mock surveys derived from different hydrodynamical simulations. Appendix D also presents results that indicate the $\nu_{X}$ parameters may not be necessary for jointly modelling the moments measurements across redshift. Bigwood et al. (2024), who use the baryonification model of Giri & Schneider (2021) to analyse data from DES and the Atacama Cosmology Telescope (ACT), also find qualitatively similar behavior to us for constraints on their mass-dependence parameter $\mu$ (see their Table B2).
@@ 6 Conclusions
The baryonification method has rapidly become a important tool in characterizing the two-point correlation functions measured by lensing surveys. The method, however, is a general prescription on how to add baryonic fields to N-body simulations through an approximate, phenomological approach. In this work, we use this formalism to model the lensing and tSZ field. One could similarly use this prescription to model other fields that are observables of gas physics, such as the kinematic SZ effect, which has already been measured using different combinations of CMB and galaxy surveys (e.g., Mallaby-Kay et al. 2023; Hadzhiyska et al. 2024), and the X-ray luminosity, measured for a large fraction of the sky by the eROSITA survey (Bulbul et al. 2024). In fact, some of these observables have already been used to constrain baryonification models (Schneider et al. 2019, 2022; Grandis et al. 2024; Bigwood et al. 2024). Thus, we have only probed some aspects of the full potential provided by the baryonification formalism.
@@ Acknowledgements
We thank Marco Gatti and Yuuki Omori for discussions during early stages of this work, Srini Raghunathan for guidance on the tSZ forward models used in Raghunathan et al. (2022a, b), and Kayla Kornoelje for details on the linear combination method used by SPT. We also thank the referee for their many helpful comments that greatly enhanced our discussion and results. The conception and pursuit of this work arose from many conversations between DA and Eric Baxter. The authors are grateful to Eric for his collaboration and friendship over the many years, and for his efforts in making our research communities vibrant avenues for discussion and discovery. He will be missed.
@@
The baryonification pipeline is publicly available at https://github.com/DhayaaAnbajagane/BaryonForge.
@@ A.1 Weak lensing
Survey mask & mass map construction. The noisy shear field — which is the sum of the true shear fields and the shape noise fields — is then masked according to the survey footprint. For DES Y6 we use the provided survey mask in the year 3 data release (Sevilla-Noarbe et al. 2021). For LSST Y10, we divide the full-sky into three equal-area cutouts of roughly 14,000 deg2 each, which is the expected area coverage (The LSST Dark Energy Science Collaboration et al. 2018, see Appendix C1). The noisy shear maps are converted to convergence using the relation in Equation 41. We only use the resulting E-mode field, $\kappa_{E}$, for our analyses. This follows the same procedures used in Chang et al. (2018); Jeffrey et al. (2021).
@@ C.2 Sensitivity to non-thermal pressure models
Figure 16 compares this model to other alternatives, particularly the predictions of Green et al. (2020). The latter model is an analytic prediction (see their Figure 1 for a summary of the model) that depends only on the peak height of the halo, and is validated using a set of hydrodynamic simulations from Nelson et al. (2014). For comparison, we plot thermal gas pressure profiles, using either Green et al. (2020) or our prescription in Section 2.3 to estimate the non-thermal pressure. In the latter approach, predictions are made for fiducial parameter values listed in Table 1. The phenomenological model, based on Shaw et al. (2010), qualitatively matches the form predicted by Green et al. (2020), which in turn is a advanced version of previous non-thermal models, such as that of Nelson et al. (2014).
@@ Appendix D Validation at higher redshifts
Appendix E Impact of parameter set of Fisher information
A number of previous works have constrained the baryonification parameters using data (e.g., Chen et al. 2023; Aricò et al. 2023; Grandis et al. 2023; Bigwood et al. 2024), and they span a variety of choices in the parameters being varied. For example Chen et al. (2023) vary a single parameter, $M_{c}$, whereas Aricò et al. (2023) varies six parameters. Additionally, Giri & Schneider (2021) explored the minimal set of parameters needed for baryonification to still predict the matter power spectrum from different simulations.
Here, we show our Fisher constraints on cosmology using different sets of baryonification parameters. Ordered by increasing complexity and number of parameters, we consider: varying no baryonification parameters, varying only $M_{c},\theta_{\rm ej}$ (base), varying the two mass-dependence parameters $\mu_{X}$, the two redshift dependence parameters $\nu_{X}$, and varying the four remaining parameters. See Table 1 for a complete list of all parameters we vary. Figure 19 shows the results of this lensing plus tSZ analysis. As expected, varying all parameters degrades cosmology constraints by roughly around $60\%$. Table 5 also lists the constraints for cosmology, as well as $M_{c},\theta_{\rm ej}$, alongside the number of baryonification parameters varied per analysis.
Figure 19: Constraints on two cosmology parameters as a function of the number of baryonifciation parameters varied in the analysis. We forecast for a DES Y6 x SPT 3G analysis, using the 2nd and 3rd moments of the fields. The constraints, and the number of baryonification parameters varied in each, are listed in Table 5. The inclusion of all 11 baryonification parameters leads to a $\approx 60\%$ degradation in cosmology constraints.
Analysis $\Omega_{\rm m}$ $\sigma_{8}$ $M_{c}$ $\theta_{\rm ej}$
Cosmo only (0) 0.005 0.009
Base (2) 0.006 0.011 0.102 0.174
Base + $M$-dep (4) 0.007 0.014 0.132 0.213
Base + $M$-dep + $z$-dep (6) 0.008 0.015 0.316 0.217
All (10) 0.009 0.015 0.554 0.478
Table 5: Constraints on two cosmology parameters, and the two key baryonification parameters (for lensing) as a function of the number of baryonifciation parameters varied in the analysis (also denoted with the brackets). We forecast for a DES Y6 x SPT 3G analysis, using the 2nd and 3rd moments of the fields.
Appendix F Computational details